### 17 Jan 2024 v1

# Yantao Linormal-$\diamondsuit$   Jianbing Zhangnormal-$\diamondsuit$   Zhiyong Wunormal-$\heartsuit$

Abstract
Graphical User Interface (GUI) agents are designed to automate complex tasks on digital devices, such as smartphones and desktops. Most existing GUI agents interact with the environment through extracted structured data, which can be notably lengthy (e.g., HTML) and occasionally inaccessible (e.g., on desktops). To alleviate this issue, we propose a visual GUI agent – SeeClick, which only relies on screenshots for task automation. In our preliminary study, we have discovered a key challenge in developing visual GUI agents: GUI grounding – the capacity to accurately locate screen elements based on instructions. To tackle this challenge, we propose to enhance SeeClick with GUI grounding pre-training and devise a method to automate the curation of GUI grounding data. Along with the efforts above, we have also created ScreenSpot, the first realistic GUI grounding dataset that encompasses mobile, desktop, and web environments. After pre-training, SeeClick demonstrates significant improvement in ScreenSpot over various baselines. Moreover, comprehensive evaluations on three widely used benchmarks consistently support our finding that advancements in GUI grounding directly correlate with enhanced performance in downstream GUI agent tasks.  [^1] The model, data and code are available at https://github.com/njucckevin/SeeClick.
1 Introduction
Developing autonomous agents to assist humans on computing devices has been a persistent goal for artificial intelligence (Shi et al., 2017; Li et al., 2020a; Furuta et al., 2023; Zhou et al., 2023). These Graphical User Interface (GUI) agent systems aim to mimic human interactions in solving complex tasks, thereby enhancing efficiency and reducing manual effort, with examples like Siri and Copilot. Recent advances in Large Language Models (LLMs) such as GPT-4 (OpenAI, 2023) and LLaMA (Touvron et al., 2023) have significantly propelled the evolution of GUI agents (Gur et al., 2023a; Zhou et al., 2023). These agents interact with the environment through extracted structured texts, e.g., HTML from web pages, then elicit LLM for planning, reasoning, and action gereration (Kim et al., 2023; Zheng et al., 2023).
However, GUI agents depend on structured text face three inherent limitations: (1) Structured text is not always accessible, especially for iOS or desktop applications where acquiring such information is challenging (Shaw et al., 2023); (2) The verbose nature of structured text serves as an inefficient context for LLMs, while also omitting crucial information such as layout, images, and icons (Deng et al., 2023); (3) The variety of structured text - including HTML, DOM, and Android VH - necessitates the curation of task-specific observation and action spaces (Kim et al., 2023; Zhou et al., 2023). These entrenched deficiencies in text-based approaches call for an alternative solution.
In this paper, we propose a visual GUI agent built on Large Vision-Language Models (LVLMs) - SeeClick. Inspired by human interaction with GUIs, SeeClick performs low-level actions like clicking or typing directly by observing interface screenshots. This methodology bypasses the need for interacting with cumbersome structured text, empowering SeeClick as a universal visual agent suitable for various GUI platforms. Building such visual agents involves a foundational challenge: GUI grounding - the capacity to accurately locate screen elements based on instructions, which is absent in current LVLMs. To tackle this challenge, SeeClick enhances LVLM with a GUI grounding pre-training strategy. We devise a method to automate the curation of web grounding data and adapt public mobile UI datasets to obtain mobile grounding data. SeeClick employs the above-curated dataset for continual pre-training of the LVLM, enabling it to accurately locate elements such as text, widgets, and icons in various GUI environments.
Given GUI grounding is a fundamental yet underexplored capacity for GUI agents, we created ScreenSpot, the first realistic GUI grounding evaluation benchmark across various GUI platforms. ScreenSpot contains over 600 screenshots and 1200 instructions from iOS, Android, macOS, Windows, and web environments, and specifically includes both text-based elements and a variety of widgets and icons. Evaluation results confirm SeeClick’s superiority over current LVLMs, validating the effectiveness of GUI grounding pre-training.
Finally, we adapted SeeClick to three downstream agent tasks: MiniWob (Shi et al., 2017), AITW (Rawles et al., 2023), and Mind2Web (Deng et al., 2023). As a purely visual-based agent, SeeClick achieves impressive performance across three tasks. Notably, SeeClick outperforms the visual baseline Pix2Act with only 0.3% training data on MiniWob. Moreover, experimental results on three tasks consistently support our finding that improvement in GUI grounding directly correlates with enhanced agent task performance.
This paper makes the following contributions:
  • We develop a unified visual GUI agent SeeClick, which directly performs clicking and typing actions based on interface screenshots across diverse GUI platforms.
  • We prospectively explore GUI grounding for visual GUI agents, and enhanced SeeClick with our proposed GUI grounding pre-training strategy.
  • We create a realistic GUI grounding benchmark ScreenSpot, encompassing more than 1200 instructions from various GUI platforms.
  • Experimental results on ScreenSpot and three agent tasks demonstrate that enhancing agents’ grounding capacity is key to improving performance in downstream agent tasks.
2 Related work
2.1 Autonomous GUI Navigation
Autonomous GUI navigation requires intelligent agents to tackle complex tasks in various GUIs. Early research investigated training agents by reinforcement learning (Liu et al., 2018; Gur et al., 2018) on the classic web environment MiniWob (Shi et al., 2017). With the recent advancement of LLMs (OpenAI, 2023; Touvron et al., 2023; Xu et al., 2023; Sun et al., 2023; Kim et al., 2023, inter alia). LLM-centric GUI agents have become the dominant paradigm. A line of works focused on prompting closed-source LLMs like ChatGPT and GPT-4, leveraging their powerful reasoning and planning abilities for web tasks, e.g., via in-context learning (Zheng et al., 2023) and self-refine (Kim et al., 2023). Other research explored training LLMs as specialized GUI agents (Gur et al., 2023b; Furuta et al., 2023). Deng et al. (2023) devised a two-stage method for identifying target elements within intricate HTML. Recently, Gur et al. (2023a) designed a method for interacting with websites via synthesizing programs.
However, constrained by the inherent limitation of LLMs to only process text and the complexity of GUIs, converting interfaces into structured text is cumbersome and challenging to standardize across various GUI environments. To address this, several recent efforts have attempted vision-based GUI navigation using GPT-4V (Yan et al., 2023; Gao et al., 2023; Yang et al., 2023). Recent studies explored vision-based UI Agents for web (Shaw et al., 2023) or mobile (Zhan and Zhang, 2023). In this paper, we construct a universal visual GUI agent SeeClick by customizing LVLMs, capable of operating across various GUI platforms.
2.2 Large Vision-Language Models
Recently, researchers have invested tremendous effort in constructing LVLMs capable of jointly processing image and text (Liu et al., 2023a; Zhu et al., 2023; Ye et al., 2023b; Li et al., 2023; Zhang et al., 2023a). LVLMs integrate visual encoding modules, like ViT (Dosovitskiy et al., 2020), with LLMs such as LLaMA (Touvron et al., 2023), inheriting LLMs’ language and reasoning abilities to perform diverse downstream vision-language tasks.
A series of studies focused on the grounding capabilities of LVLMs (Wang et al., 2023; Bai et al., 2023; Chen et al., 2023a), such as providing bounding boxes for objects while generating responses (Chen et al., 2023b; Peng et al., 2023). Nonetheless, these efforts primarily addressed natural images and did not explore GUI contexts. This paper emphasizes that GUI grounding is pivotal for constructing visual GUI agents and concentrates on exploring LVLMs’ potential in GUI environments.
Figure 1: Overview of our universal visual GUI agent SeeClick. (a) depicts the framework of SeeClick and GUI grounding pre-training. (b) provides examples of ScreenSpot across various GUIs and types of instructions. (c) displays the real-world application of SeeClick on downstream web agent task.
2.3 Mobile UI Understanding
As a vital aspect of daily life, mobile UI understanding has garnered widespread attention from researchers. One category of work explored multimodal tasks within UI contexts (Li et al., 2021; Bai et al., 2021; He et al., 2021), such as widget captioning (Li et al., 2020b), screen summarization (Wang et al., 2021), and UI object detection (Chen et al., 2020). Several studies explored grounding elements in mobile UI based on instructions (Li et al., 2020a; Burns et al., 2022; Li et al., 2021; Li and Li, 2022), serving as a foundation for mobile UI automation. Most mentioned studies rely on Android’s View Hierarchy to represent interfaces, while (Li and Li, 2022) use Regions of Interest and (Zhang et al., 2023b) use screenshots. Moreover, most efforts employed RICO (Deka et al., 2017) for research, an Android apps dataset collected in 2017. This work introduces an up-to-date, realistic grounding benchmark encompassing various GUI platforms, and underscores grounding as a foundation for visual GUI agents.
3 Approach
In our preliminary study, we have discovered a key challenge in developing visual GUI agents: GUI grounding, the capacity to locate screen elements based on instructions. Recent advancements in LVLMs have claimed visual grounding capability on natural images (Chen et al., 2023b; Peng et al., 2023; Wang et al., 2023), which is able to point out objects based on natural language description. However, GUI screenshots significantly differ from natural images, typically containing dense text and numerous icons and widgets. These differences impair existing LVLMs’ grounding performance in GUI contexts (Bai et al., 2023; Chen et al., 2023a), limiting their potential as visual GUI agents.
This paper seeks to endow LVLMs with GUI grounding proficiency, aiming to construct a universal visual GUI agent capable of executing instructions that only rely on screenshots. The overview of the proposed LVLM SeeClick is presented in Figure 1(a). Next, we introduce the birth of SeeClick, including the formalization of GUI grounding task, the construction of continual pre-training data, and training details.
3.1 GUI grounding for LVLMs
As GUI grounding is the core capability of SeeClick, we first elucidate how to train LVLM for language generation to perform grounding tasks. Given an interface screenshot $s$ and a collection of its elements $\{(x_{i},y_{i})|_{i}\}$, where $x_{i}$ denotes the textual description of the $i$-th element and $y_{i}$ indicates the element’s location (represented as a bounding box or point). As depicted in Figure 1(a), LVLM predicts the location of the element $y$ based on the interface screenshot $s$ and its textual description $x$, i.e. calculating $p(y|s,x)$.
A potential challenge is how LVLMs predict numerical coordinates in a language generation format. Traditional methods (Chen et al., 2021; Wang et al., 2023; Shaw et al., 2023) divided the image into 1000 bins, and creating a new 1,000-token vocabulary $\{<p0>,<p1>,...,<p999>\}$ to represent the x and y coordinates. In this work, we adopt a more intuitive manner used in LVLMs (Chen et al., 2023b; Bai et al., 2023), treating numerical values as natural language without any additional tokenization or pre-/post-processing. For instance, in Figure 1(a), for a smartphone screenshot and the instruction “View the new album of Jony J”, we craft a user query prompt: “In the UI, where should I click if I want to <instruction>?”. Subsequently, we normally compute the cross-entropy loss between the model output and the groundtruth “click (0.49, 0.40)” to optimize the LVLM.
Figure 2: Example of two types of elements automatically collected from the webpage.
3.2 Data Construction
We primarily train SeeClick using three collections of data: web UI data crawled from the internet, mobile UI data reorganized from public datasets, and general vision-language instruction-following data.
Web Data. Web UIs, featuring a variety of layouts and design styles across websites, are ideal for training LVLMs to learn universal recognition, understanding, and grounding capabilities across different GUIs. We collected approximately 300k web pages from the latest Common Crawl repository  [^2] https://commoncrawl.org/ to serve as our training data for web UI. For each webpage, we collect two types of elements from the HTML code: (1) elements that display visible text content; and (2) elements with a special “title” attribute that display descriptive text when hovering. This method ensures that we gather a series of interactable elements and their corresponding text instructions, while encompassing a wide range of texts and icons. Figure 2 provides an example of a web page, its associated simplified HTML code and extracted two type elements. Next, we treat the text or “title” content of each element as the instruction, constructing the GUI grounding data with screenshot $s$, instruction $x$, and the element’s location $y$. We also include an inverse task, similar to web OCR, where given an interface screenshot $s$ and the location $y$ of an element, the model predicts the text description $x$ of the element.
Mobile Data. For mobile UI, we include three types of data: widget captioning, mobile UI grounding, and mobile UI summarization. Widget Captioning aims to automatically generate language descriptions for mobile UI elements; for example, producing the description “play music” for the play button on a music player interface. We utilize the training split of the dataset provided in (Li et al., 2020b), containing nearly 20k screenshots, 40k widgets and 100k descriptions. We derive mobile UI grounding data by inverting the widget captioning task, treating language descriptions as instructions and predicting widget locations from screenshots and instructions. To improve element diversity, we also incorporate the automatically generated elements and instructions from RICO (Li et al., 2020a). The mobile UI grounding data involve diverse elements and instructions, aids SeeClick in generalizing grounding capabilities to various GUI scenarios. We finally include mobile UI summarization data (Wang et al., 2021) to enhance overall interface comprehension.
General Data. To maintain LVLM’s general capacities on natural images and its planning and reasoning abilities for downstream agent tasks, we incorporate general vision-language data for training SeeClick. We utilize the 158k instruction-following samples collected with GPT-4 from LLaVA (Liu et al., 2023a), covering conversation, detailed description, and complex reasoning.
Finally, we mix the data above and craft 30 task-specific prompts for each added GUI task, resulting in a 1M dataset to train SeeClick.
3.3 Training Details
We built SeeClick through continual pre-training on a recent advanced LVLM, Qwen-VL (Bai et al., 2023). Qwen-VL inherently possesses grounding capabilities and a higher resolution (448*448), which has been proven beneficial for understanding text-rich images (Lee et al., 2023; Ye et al., 2023a). We train Qwen-VL on the constructed dataset (Section 3.2) for about 10k steps (around 1 epoch) to obtain our GUI base model SeeClick. During training, we employ LoRA (Hu et al., 2021) to fine-tune both the visual encoder and LLM. We adopt AdamW as the optimizer and use a cosine annealing scheduler with an init learning rate of 3e-5 and a global batch size of 64. All training takes around 24 hours on 8 NVIDIA A100 GPUs.
Figure 3: Statistic of our proposed GUI grounding benchmark ScreenSpot. The left illustrates the diverse GUI environments included. The right displays the types of elements within each GUI category.
Table 1: Results of different LVLMs on ScreenSpot. The best results in each column are bold. Benefit from efficient GUI grounding pre-training, SeeClick significantly enhanced LVLMs’ ability to locate GUI elements following instructions, and surpassed the strong baseline CogAgent with a smaller model size.
LVLMs Model Size GUI Specific Mobile Desktop Web Average
Text Icon/Widget Text Icon/Widget Text Icon/Widget
MiniGPT-v2 7B 7.6% 5.2% 4.5% 2.9% 5.5% 0.0% 4.3%
Qwen-VL 9.6B 9.4% 4.3% 6.4% 5.3% 0.0% 0.0% 4.2%
Fuyu 8B 40.6% 1.6% 33.6% 6.7% 48.4% 2.9% 22.3%
CogAgent 18B 66.5% 26.7% 73.7% 19.3% 78.0% 21.4% 47.6%
SeeClick 9.6B 75.9% 48.7% 72.7% 26.4% 69.2% 21.4% 52.4%
4 GUI Grounding Benchmark: ScreenSpot
We recognize GUI grounding proficiency as essential for constructing a universal visual GUI agent. However, due to the constrained capabilities of earlier vision-language models, previous research has paid limited attention to GUI grounding, with the few existing studies (Li et al., 2021; Li and Li, 2022; Zhang et al., 2023b) largely confined to Android environments.
To address this research gap, we introduce ScreenSpot, an up-to-date, realistic grounding evaluation benchmark encompassing various GUI platforms. It is designed to assess vision-language models’ ability to locate elements based on human instructions (Figure 1(b) provides some examples). ScreenSpot has two distinctive features: (1) Various GUI platforms. It includes over 600 interface screenshots from mobile (iOS, Android), desktop (macOS, Windows), and web platforms, along with more than 1200 instructions and corresponding actionable elements (as shown in Figure 3 left); (2) Icons/Widgets. Considering that locating icons or widgets might be more challenging than text elements on GUIs, we include a substantial number of icon/widget type elements in each GUI (as depicted in Figure 3 right). See Appendix A for more examples.
To evaluate vision-language models’ effectiveness in real-world scenarios, ScreenSpot is carefully curated to ensure the samples in ScreenSpot are novel and not included in existing training resources. We recruited experienced annotators to collect new GUI interfaces and label instructions along with the bounding boxes for actionable elements. For mobile and desktop, annotators were asked to select commonly used apps and operations; for web pages, we chose several types of websites (development, shopping, forum, and tools) from the recent web environment WebArena (Zhou et al., 2023) for annotation.
5 Experiments
In this section, we first evaluate the GUI grounding capabilities of representative LVLMs and our proposed SeeClick on ScreenSpot. Subsequently, we adapt SeeClick to three downstream GUI agent tasks, exploring and analyzing the performance and potential of vision-based agents in various GUIs.
5.1 GUI Grounding on ScreenSpot
As the foundation of visual GUI agents, GUI grounding has not received extensive attention in current LVLMs evaluations (Liu et al., 2023b; Yu et al., 2023). Therefore, we compare various LVLMs on our specially constructed GUI grounding benchmark ScreenSpot, to examine their potential for GUI automation.
Compared LVLMs & Evaluation. We primarily evaluated two types of LVLMs: (1) versatile general LVLMs capable of tasks such as dialogue, recognition and grounding, including MiniGPT-v2 (Chen et al., 2023a) and Qwen-VL (Bai et al., 2023); (2) recent LVLMs specifically designed for GUI tasks, including Fuyu (Bavishi et al., 2023), CogAgent (Hong et al., 2023) and our SeeClick.
Considering that GUI agents require clicking on the correct position, we calculate the click accuracy as the metric, defined as the proportion of test samples where the model’ predicted location falls in the groundtruth element bounding box (Zhang et al., 2018; Li et al., 2022; Zhang et al., 2023b). We take the models’ generated point or the center of the generated bounding box as the predicted click location.
Results. As shown in Table 1, while general LVLMs have excelled in natural image grounding (e.g., RefCOCO (Kazemzadeh et al., 2014)), their GUI grounding performance on ScreenSpot is poor due to the significant differences between GUIs and natural images.
Accordingly, LVLMs trained specifically on GUI data have significant improvements. Our SeeClick achieved the best average performances across three GUI platforms and two types of elements, even with fewer model parameters and GUI-related training data than CogAgent. This demonstrates the efficiency of our GUI grounding pre-training; with the rich UI elements and diverse instructions collected from the web and mobile, SeeClick quickly learns to understand human instructions and locate elements, even in completely unseen scenarios like iOS and desktop. Notably, all models struggle with locating icons/widgets, highlighting the difficulty of identifying and grounding non-text elements on GUIs, and showcasing the unique challenges posed by ScreenSpot.
5.2 Visual GUI Agent Tasks
This section explores the capabilities of SeeClick when migrated to three downstream GUI agent tasks: MiniWob, AITW, and Mind2Web. Across these tasks, SeeClick emulates human behavior, determining the next action solely by observing interface screenshots, without needing any additional GUI information. On all three tasks, we trained SeeClick on the respective training splits and tested it on the test sets. Appendix B provides examples of SeeClick’s interactions in these tasks.
Figure 4: Comparison of SeeClick and LVLM baseline Qwen-VL on MiniWob. Tasks with yellow shadows indicate those with dynamic webpage layouts and element positions (closer to the real-world application of GUI agents, see more details in Section B.1). SeeClick outperformed Qwen-VL in most tasks, especially in those with dynamic elements, highlighting the effectiveness of our GUI grounding pre-training.
5.2.1 MiniWob
MiniWob Shi et al. (2017) comprises about 100 types of web automation tasks, where the agent is asked to interact with a simplified web environment to accomplish human instructions. Each task can templatize random variants and corresponding instructions controlled by a random seed, creating up to billions of possible task instances.
Existing open-source training data often lacks corresponding interface screenshots (Furuta et al., 2023). Therefore, we rollout 50 successful episodes using an LLM strategy for each task in (Zheng et al., 2023), resulting in a 2.8K episodes dataset to train SeeClick.
Compared Methods & Evaluation. We compared SeeClick with a range of offline training methods. Among these, the state-of-the-art method WebGUM (Furuta et al., 2023) uses screenshots as auxiliary information but still selects HTML elements as actions. Pix2Act (Shaw et al., 2023) is the only prior vision-based approach, trained with extensive data to perform clicking and typing actions on MiniWob. To verify the effectiveness of GUI grounding pre-training, we also report the results of the LVLM baseline Qwen-VL when trained with the same 2.8K dataset.
Due to the variance in evaluation task sets among different methods (Liu et al., 2018; Furuta et al., 2023; Shaw et al., 2023), for fairness, we report performance in two groups based on the overlapping MiniWob tasks with their approaches. We compute the success rate over 50 random seeds for each task, and then compute the mean over all MiniWob tasks as the final score.
Table 2: Average scores of different methods on MiniWob. The best results in each setting are bold. Methods achieving the highest performance with limited data are underlined. SeeClick outperforms a range of offline training methods as a purely vision-based model.
Methods Modality Dataset Score
Compared with text-based models over 45 tasks
CC-Net (SL) DOM+Image 2.4M 35.6%
WebN-T5 HTML 12K 55.2%
MM-WebN-T5 HTML+Image 347K 63.4%
WebGUM HTML+Image 2.8K 65.5%
WebGUM HTML+Image 347K 86.1%
SeeClick Image 2.8K 73.0%
Compared with vision-based models over 35 tasks
CC-Net (SL) Image 2.4M 23.4%
Pix2Act Image 1.3M 64.6%
Qwen-VL Image 2.8K 48.4%
SeeClick Image 2.8K 68.0%
Table 3: Average scores of different methods on AITW. ClickAcc calculates the accuracy of click operation. The best results in each column are bold. SeeClick exhibits the best performance among both API-based LLMs (top row) and trained LVLMs (bottom row).
Methods Modality General Install GoogleApps Single WebShopping Overall ClickAcc
ChatGPT-CoT Text 5.9 4.4 10.5 9.4 8.4 7.7 -
PaLM2-CoT Text - - - - - 39.6 -
GPT-4V Image 41.7 42.6 49.8 72.8 45.7 50.5 -
Qwen-VL Image 53.3 61.2 56.3 61.7 52.0 56.9 60.1
SeeClick Image 56.0 64.5 57.7 63.6 57.3 59.8 68.4
Results. As depicted in Table 2, our purely vision-based approach SeeClick surpassed a series of HTML-based and image-based baselines with only a fraction of the training data. Notably, with an equivalent amount of training data (2.8K), SeeClick outperformed the previous offline sota WebGUM, which uses both HTML and screenshots as input. In comparison with visual approaches, thanks to the powerful reasoning and planning abilities of the LVLM and the localization skill imparted by our GUI grounding pre-training, SeeClick exceeded the sota visual method Pix2Act, using less than $0.3\%$ training data.
Furthermore, SeeClick significantly surpassed the LVLM baseline Qwen-VL by 20 percentage points, verifying that enhancing GUI grounding capabilities plays a crucial role in improving LVLMs’ performance on MiniWob tasks. To analyze in detail, we provide task-level comparisons between SeeClick and Qwen-VL on 35 MiniWob tasks in Figure 4. SeeClick shows notable improvements over Qwen-VL in most tasks, particularly in those where interface layouts and element positions dynamically change. This confirms our hypothesis that general LVLMs struggle with accurately clicking on interfaces, and our GUI grounding pre-training significantly enhances this capability, leading to better performance.
5.2.2 AITW
We evaluate the effectiveness of SeeClick in smartphone environments with a recently proposed Android UI automation dataset Android In The Wild (AITW). It encompasses 30k distinct task instructions and corresponding 715k human-annotated operation episodes. Previous approaches (Rawles et al., 2023; Zhan and Zhang, 2023) split AITW episode-wise into train/val/test (80/10/10%). However, with an average of over 20 episodes per instruction in AITW, there is a significant risk of overfitting due to: (1) extensive repetition in training trajectories; (2) instructions in the test set have appeared during training. In this work, we propose to split the training and testing sets by instruction. Specifically, we selected 545/688/306/700/700 instructions from General/Install/GoogleApps/Single/WebShopping respectively and retained only one annotated trajectory for each instruction. Then for each scenario, we selected 80% for training and the remaining for testing. This setting better reflects the performance of the trained agent in the real-world application.
Compared Methods & Evaluation. We compare SeeClick with two types of baselines: (1) API-based LLMs employing in-context learning or Chain-of-Thought (CoT), such as ChatGPT-CoT (Zhan and Zhang, 2023), PaLM2-CoT (Rawles et al., 2023) and the latest GPT-4V (Yan et al., 2023); (2) our own trained LVLM baseline Qwen-VL. We follow Rawles et al. (2023) to adopt the screen-wise action matching score as the main metrics, and compute the click accuracy (ClickAcc), which calculates the accuracy when both the reference and prediction are click actions.
Results. As illustrated in Table 3, SeeClick achieved the best average performance among both API-based LLMs and trained LVLMs. Specifically, SeeClick exhibited an 8 percentage point improvement in click accuracy over the LVLM baseline Qwen-VL, substantiating the hypothesis that GUI grounding significantly augments the precision of click actions in visual GUI agents, thereby improving their overall task performance.
5.2.3 Mind2Web
To assess SeeClick’s capabilities in web environments, we utilize the recently introduced Mini2Web dataset (Deng et al., 2023). Mind2Web comprises over 2000 open-ended tasks collected from 137 real websites, each with a high-level instruction and a human action trajectory sequence involving clicking, selecting and typing. Mind2Web was originally designed to evaluate text-based web agents, focusing on identifying actionable elements by selecting them from simplified HTML documents. This work explores visual web agents that predict click positions directly from interface screenshots. For this purpose, we augmented Mind2Web’s trajectory data with interface screenshots and bounding boxes for action elements at each step, sourced from their raw dump data (Deng et al., 2023). To the best of our knowledge, this is the first attempt of web agents relying solely on screenshots as inputs for navigating real websites.
Table 4: Comparsion of different methods on Mind2Web. The best results in each column are bold. Improvements of SeeClick over the LVLM baseline are underline. The GUI grounding pre-training nearly doubled the step success rate of SeeClick on real websites.
Methods w/o HTML Cross-Task Cross-Website Cross-Domain
Ele.Acc Op.F1 Step SR Ele.Acc Op.F1 Step SR Ele.Acc Op.F1 Step SR
MindAct (gen) 20.2 52.0 17.5 13.9 44.7 11.0 14.2 44.7 11.9
MindAct 55.1 75.7 52.0 42.0 65.2 38.9 42.1 66.5 39.6
GPT-3.5 20.3 56.6 17.4 19.3 48.8 16.2 21.6 52.8 18.6
GPT-4 41.6 60.6 36.2 35.8 51.1 30.1 37.1 46.5 26.4
Qwen-VL 14.9 86.2 12.6 12.1 83.7 10.1 9.6 84.3 8.0
SeeClick 26.3 86.2 23.7 21.9 82.9 18.8 22.1 84.1 20.2
Compared Methods & Evaluation. We compare SeeClick with a range of text-based web agents (Deng et al., 2023) and our LVLM baseline Qwen-VL. Mind2Act employs a two-stage method, where a small LM first generates candidate elements from raw HTML, and then a large LM selects the target element using multi-choice QA. Mind2Act (gen) modifies the QA module to directly generate the target element instead. GPT-3.5 and GPT-4 adopt the same multiple-choice QA formulation as MindAct, and include three demonstration examples for in-context learning.
We calculate element accuracy (Ele.Acc), Operation F1 (Op.F1) and step success rate (Step SR). For vision-based methods, a prediction is considered correct if the predicted coordinate falls in the target element’s bounding box. All other settings are following (Deng et al., 2023).
Results. As displayed in Table 4, SeeClick nearly doubled the element selection accuracy and step success rate compared to the LVLM baseline Qwen-VL. This indicates that improvements in GUI grounding directly correlate with enhanced performance in downstream web agent tasks. Furthermore, text-based methods often have lower overall Op.F1 due to nearly 20% groundtruth elements being incorrectly filtered out in the candidate generation stage, which leads to a zero Op.F1.
Although SeeClick can operate without extra HTML information, its performance lags behind state-of-the-art HTML-based methods, as predicting click locations on web pages is much more difficult than choosing from HTML candidates. This underscores the challenge of grounding in complex interfaces, suggesting substantial room for improvements toward a visual GUI agent that is applicable in real-world scenarios.
6 Conclusion
In this paper, we introduce a visual GUI agent - SeeClick, which only relies on screenshots for GUI task automation. We found a key challenge in developing such visual GUI agents: GUI grounding - the capacity to accurately locate screen elements based on human instructions. To address this challenge, we propose to enhance SeeClick via GUI grounding pre-training, and devise methods to automate the curation of GUI grounding data from web and mobile. For benchmarking the progress in GUI grounding, we created ScreenSpot, the first realistic evaluation dataset encompassing mobile, desktop, and web platforms. Results on ScreenSpot demonstrate a significant improvement of SeeClick over LVLM baselines. Moreover, comprehensive evaluations across three GUI automation tasks consistently support our finding that advancements in GUI grounding directly correlated with improved performance in downstream agent tasks.
Appendix A Examples of ScreenSpot
Figure 5 provides more examples of ScreenSpot, which contains a variety of common GUI scenarios for mobile, desktop, and web platforms.
Appendix B Case Study
B.1 MiniWob
Figure 6(a) illustrates the difference between static and dynamic layout tasks. Static layout tasks have fixed element positions during training and testing, while dynamic layout tasks display varying interfaces and element positions. Figure 6(b) provides examples of SeeClick’s interaction with MiniWob. SeeClick relies solely on the interface screenshot for arithmetic, reasoning, etc.
B.2 AITW
Figure 7 provides an example of SeeClick when completing an AITW task. Predictions marked in red below indicate that they were computed as incorrect in AITW. However, some errors occur because the current step’s answer is not unique. For example in step 5, the model’s predicted input "DuckDuckGo Privacy Browser" is also a potentially correct action.
B.3 Mind2Web
Figure 8 shows several examples of SeeClick on the real-world website benchmark Mind2Web. SeeClick can comprehend instructions and click on the correct elements within complex interfaces.
Figure 5: More examples of GUI grounding benchmark ScreenSpot.
Figure 6: Example episodes of SeeClick on MiniWob. The model’s prediction output is below the screenshot, with action_type 4 indicating clicking and action_type 3 indicating typing.
Figure 7: Example episodes of SeeClick on AITW. The model’s prediction output is below the screenshot, with action_type 4 indicating clicking, action_type 3 indicating typing and action_type 6 indicating PRESS HOME. For failed steps, we mark the predictions in red and provide the groundtruth action in green.
Figure 8: Example episodes of SeeClick on Mind2Web. The model’s prediction output is below the screenshot, with action_type 4 indicating clicking and action_type 3 indicating typing. For failed steps, we mark the predictions in red and provide the groundtruth elements with the green bounding box.